OpenStudy (anonymous):

@johnweldon1993 @jmoney132 @justin_lane47 Factor the polynomial completely: 3x2 - 4x + 21 A) (3x + 7)(x + 3) B) (x - 7)(3x - 3) C) (x - 3)(3x - 7) D) cannot be factored

OpenStudy (anonymous):

mmm math isnt my strong suit hun sorry

OpenStudy (anonymous):

what are you good at then

OpenStudy (maddy1251):

I can help

OpenStudy (maddy1251):

Since this is MC, its actually gonna be easier on your part. Do you know how to foil

OpenStudy (anonymous):

yes but i dont like it

OpenStudy (maddy1251):

I know, anything like that isn't fun. However, we can FOIL your answer choices to see if any of them match up with 3x^2-4x+21

OpenStudy (anonymous):

my only weak one is the algebra area

OpenStudy (maddy1251):

I got this @hammerhelm :-)

OpenStudy (maddy1251):

A) (3x + 7)(x + 3) = 3x^2+9x+7x+21 B) (x - 7)(3x - 3) 3x^2-3x-21x +21 C) (x - 3)(3x - 7) 3x^2-7x-9z+21. D) cannot be factored

OpenStudy (anonymous):

alright thanks girl ttyl i got another question asked by someone

OpenStudy (maddy1251):

3x2 - 4x + 21 This is what we are looking for.

OpenStudy (maddy1251):

(3x + 7)(x + 3) = 3x^2+9x+7x+21 =3x^2+16x+21. Nope. B) (x - 7)(3x - 3) 3x^2-3x-21x +21 3x^2-24x+21. Nope. C) (x - 3)(3x - 7) 3x^2-7x-9z+21. 3x^2-16x+21. Nope.

OpenStudy (maddy1251):

When I went to Tigermath.com for the facor form of this number, there was a radical in it. That is a little bit calculus to explain that. If you go here, you can see. http://www.tiger-algebra.com/drill/3x~2-4x_21=0/

OpenStudy (maddy1251):

It can be factored, just the correct answer choice is not listed.